OpenStudy (anonymous):

A coin is flipped four times in a row. What is the probability that the result includes exactly one tails?

OpenStudy (tkhunny):

Expand this... \((p+q)^{4}\) ...and ponder

OpenStudy (anonymous):

What is the p and the q?

OpenStudy (tkhunny):

Maybe it's a surprise? Did you do it?

OpenStudy (anonymous):

I have no idea what exactly you're saying?

OpenStudy (tkhunny):

You're saying you've never multiplied a polynomial? (p+q)(p+q)(p+q)(p+q)

OpenStudy (tkhunny):

Please read up on the Binomial Theorem and Binomial Distribution \(p^{4} + 4p^{3}q + 6p^{2}q^{2} + 4pq^{3} + q^{4}\) Now suppose p = p(heads) and q = 1-p = p(tails)

OpenStudy (tkhunny):

btw The words "in a row" are meaningless in this problem statement.

OpenStudy (anonymous):

No it's probability. I just forgot how to figure it out. I've multiplied polynomials. I don't understand where you got p and q

OpenStudy (tkhunny):

What do you mean "no"? Are polynomials somehow illegal in probability? You are performing a Binomial Experiment. The Probability of Heads is 1/2. The Probability of Tails is 1/2. You perform the experiment 4 times. You might find it under "Binomial Distribution", n = 4, p = 1/2